OpenStudy (anonymous):

find the discriminant 2x^2-9x+7=0

OpenStudy (anonymous):

ax^ + bx + c compare with ur equation 2x^2 - 9x + 7 so a=2,b=-9 and c = 7 plug these into b^2 - 4ac and you have it

myininaya (myininaya):

\[b^2-4ac\] you determine what a, b, c are by comparing your equation to \[ax^2+bx+c=0\]

OpenStudy (anonymous):

-137?

myininaya (myininaya):

no b=-9 a=2 c=7 so we have \[(-9)^2-4(2)(7)\]

OpenStudy (anonymous):

Now I came up with 25

myininaya (myininaya):

right (-9)^2=81

OpenStudy (anonymous):

o the discriminant is 25

myininaya (myininaya):

81-8(7)=81-56=25

OpenStudy (anonymous):

Ok thanks a bunch!! I really appreciate it